What is this firm's profit if price, p = $72?

Profit = pq - 6q2 - 12

432

0

78

204
Answer #1

Answer: Option d) is correct. 204

Reason:

STEP 1: First derivative of Profit

d(Profit) / dq = p -12q

Step 2: Equate it to 0 for profit maximization

p -12q = 0

p = 12q

72 = 12q

q = 6

Thus profit = pq - 6q2 - 12

= 72*12 - 6*36 - 12

= 204